# Constraining the dispersion measure redshift relation with simulation-based inference

### v1

@@ 6.2 Variations in the LSS component
Our fiducial model consisted of a log-normal realisation of the LSS component along with a log-normal host model. Next, we change the LSS component to a Gaussian distribution and observe how the results are affected, which can be done trivially with our pipeline. To that end, the SBI pipeline, as described in Section 4.2, is applied by changing the field in GLASS. For the comparison, we reduce the resolution of the simulations to $N_{\mathrm{side}}=512$. As we have established that the high-resolution run faithfully reproduces the real posterior distribution and is a good fit to the data in the last section, reducing the resolution to $N_{\mathrm{side}}=512$ makes a quantitative comparison while also requiring less computational resources. We rerun the fiducial case at the same resolution as well to make a quantitative comparison. It should be noted that the lower resolution effectively reduces the field variance of the LSS component. In appendix A we discuss the effect of reducing $N_{\mathrm{side}}$ on the coverage tests. It can be seen that the resulting posterior estimates are slightly biased when $N_{\mathrm{side}}$ is lowered, reducing the variance of the LSS component. Since we have established unbiased estimators at high resolution and checked that the constraints and $\chi^{2}$-test are not affected, this does not hamper our analysis when comparing different models.
@@
The contours are shown in Figure A.2 and the numerical values of the means and the $1\sigma$ errors of the parameters are presented in Table 2. As can be observed, there is a significant overlap among the values. Only by observing the contour plots, we cannot distinguish the models. Hence, we rely on the $\chi^{2}$-tests, shown in Figure A.3 respectively. Even then, judging from these figures, there is no discernible difference between the Gaussian and log-normal LSS components. All the variations seem to be good fits according to the likelihood in Figure A.4. Consequently, with the current data, one cannot distinguish a Gaussian from a log-normal LSS component. A larger number of host identified FRBs is required for meaningful detection of this difference. In particular, FRBs with higher redshift would be especially important to assess the effect due to the LSS field, as it becomes more dominant as the redshift increases.
@@
6.3 Variations in the Host component
Now, we turn our focus on the host model of the DM, changing it from the fiducial log-normal to first a truncated Gaussian (t-Gaussian hereafter) and then to a Gamma distribution. The Gaussian is truncated at zero as the DM is positive by definition, hence, it is not a Gaussian mathematically. For the t-Gaussian and Gamma hosts, the prior on $\sigma$ is uniform on [0, 500], which is broad enough to capture the expected long-tail behaviour. It also implies that the Gamma distribution can resemble either a log-normal or a t-Gaussian distribution depending on the data, as all of these distributions come from the exponential family. As before, the contours, $\chi^{2}$ and likelihood are shown in Figure A.2 Figure A.3 and Figure A.4 respectively, all with $N_{\mathrm{side}}=512$.
@@ 6.3 Variations in the Host component
Looking at the best fit values, we can see that the statistical properties of the LSS field are rather sub-dominant with the current FRB sample and if a log-normal distribution is used for the host contribution. If we assume a (truncated) Gaussian host contribution, it cannot explain the high DM of FRBs at low redshifts, Hence the model responds by artificially increasing the variance of the host contribution when the LSS is chosen to be Gaussian as well. In general, however, the constraints are all consistent with each other.
@@ 7 Conclusion
In the next step, we performed SBI on the described simulations. This inference method requires no explicit likelihood and works by training an NN to learn either the posterior distribution (non-amortised) or the joint distribution of data and parameters (amortised). In this context, the shape of the likelihood is unconstrained, which is why this approach is often referred to as likelihood-free inference (though it still implicitly requires a likelihood). To demonstrate our pipeline, we applied it to 12 host-identified FRBs as a function of the cosmological amplitude of the DM-$z$ relation, $\mathcal{A}$ in Equation 7, and two host contribution parameters. For our fiducial model, we used a log-normal electron density field as well as a log-normal host model. We inferred $\mathcal{A}=0.89^{+0.31}_{-0.29}$, consistent with unity or the Planck cosmology. Similarly, the median and scale of the host distribution are $\mathrm{DM}_{\mathrm{host}}=223^{+108}_{-95}\;\mathrm{pc\;cm}^{-3}\ \mathrm{and}\ \sigma_{\mathrm{LN}}=0.76^{+0.37}_{-0.44}$. The values for these three parameters are consistent with previous findings, and we indeed found that no parameter (amplitude of the DM-$z$ relation, mean host contribution and its variance) is dominated by its prior, i.e. that there is additional information in the data and no parameter just assumes a flat posterior in the prior range. This is indicating that already 12 FRBs can inform us about the shape of the host contribution to some extent.
@@
One of the added benefits of the SBI pipeline is that we can test the Gaussian likelihood assumption frequently used in a more traditional Bayesian approach. To that regard, we investigated the shape of the likelihood, finding the expected long-tailed distribution towards high DM values, which is necessary to explain the large DMs observed at low redshifts, such as those seen in FRB 20190520B. Another important check involved examining the evolution of the likelihood with redshift. We found that the mean of the probability distribution function increases with redshift, as expected. However, we also observed a reduction in its scatter. Although this behaviour may seem counterintuitive at first, it is supported by a straightforward analytical calculation, which shows that the width of the total distribution indeed reaches a minimum at redshift $z<1$ for the parameter values assumed in this analysis.
Lastly, we explored different modelling choices for both the LSS component and the host contribution. For the LSS component, we considered both Gaussian and log-normal distributions, while for the host contribution, we evaluated log-normal, truncated Gaussian and gamma distributions. We systematically tested all possible combinations of these models and found that they consistently provided similar constraints. To save computational time, we reduced the resolution of these simulations from $N_{\mathrm{side}}=4048$ to $N_{\mathrm{side}}=512$. We found that the TARP test indicates that the learned posterior in these cases is slightly biased. Importantly, none of the models showed any indication of being a poor fit to the data, which is mainly due to the low number of FRBs available with host identification at the moment.
In conclusion, the simulations and inference pipeline we have developed integrate the precise physical and statistical properties needed to accurately infer the DM-$z$ relation, free from general assumptions about the likelihood or posterior. This approach is highly adaptable and scales effectively with an increasing number of FRBs, thanks to the implemented data compression techniques. Moreover, our simulations provide a robust foundation for investigating systematic effects in cosmological studies involving FRBs, as these can be seamlessly incorporated at the map level. Looking ahead, a key direction will be to include more FRBs without known redshifts, enabling a joint fit of the DM-$z$ properties alongside the statistical properties of the DM.
@@ Appendix A Effect of reducing $N_{\mathrm{side}}$
Figure A.1: The multivariate TARP coverages for the fiducial model (using a log-normal distribution for both the field and the host) These are computed from 1000 forward simulations. One can see that and increased resolution of the maps increases the accuracy of the final posterior estimate. This behaviour is found in all cases studied in this work.
In Figure A.1, we show the effect of reducing the resolution of the GLASS forward simulation on the TARP coverage. This is necessary to establish the baseline for the comparison of the different models for the host and the LSS contribution, which have been run at a lower resolution. The main effect of decreasing $N_{\mathrm{side}}$ is that power on small scales gets washed out. In particular, the maximum multipole properly resolved is $\ell_{\mathrm{max}}=3N_{\mathrm{side}}-1$. However, loss of power already occurs earlier. In terms of effects on the inference process, this means that the variance introduced by the LSS component decreases and therefore the scatter in the data will be larger than in the simulations. Therefore, the final constraints might be artificially tight. Figure A.1 demonstrates that the estimated posterior becomes slightly biased for decreasing $N_{\mathrm{side}}$ for the fiducial model. Since this will be true for all models and the fact that the high resolution run is unbiased and exact, we can safely do a one-to-one comparison of the different models using a lower resolution.
Figure A.2: Marginalised contour plots for different combinations of the density field and host distribution. Each panel have the same host distribution and only the density field is varied, where the blue colour is for the log-normal field and red is for the Gaussian field. The host distribution for the top left, top right and bottom panels are log-normal, t-Gaussian and Gamma respectively. The posteriors for all of these contours are evaluated with 1000 forward simulations in a 10-round TSNPE setup. The resolution of all the contours is $N_{\mathrm{side}}=512$ for comparison. The means with the $1\sigma$ errors are provided in Table 2.
@@ Appendix B Results from different fits
In this section, we present a comprehensive analysis of the different figures resulting from variations in both the host contribution and the large-scale structure (LSS) contribution. Specifically, we display the final contour plots Figure A.2, the goodness-of-fit tests in Figure A.3 and the likelihoods in Figure A.4. Note that we $N_{\mathrm{side}}=512$ in this comparison to conserve computational resources. The reduced resolution underestimates the error from cosmic variance on individual measurements, leading to artificially tighter constraints. This issue has been addressed in our fiducial run, where a higher $N_{\mathrm{side}}$ was utilised to validate the pipeline, as described in the main text. The key conclusion from these analyses is that all combinations of the model are statistically consistent with one another and provide a satisfactory fit to the data. This is primarily due to the relatively small number of FRBs considered in this study, as well as the conservative error estimates and the expansive posterior volume resulting from the use of a three-parameter model. While these effects limit the current discriminative power, they ensure robustness in our findings.
@@
Finally, for reference, we provide a list of the FRBs used in this analysis in Footnote 4.
@@
Figure A.4: Likelihoods for all the models in our analysis. The two rows represent the log-normal and Gaussian density fields and the three columns are for the log-normal, t-Gaussian and Gamma host distribution. The simulated DMs for the individual FRBs are shown with a colour bar by mapping it to the percentile values. The cyan dots are the median and the black crosses are the observed DM.

### v2

@@ 6.2.1 Variations in the LSS component
Our fiducial model consisted of a log-normal realisation of the LSS component along with a log-normal host model. Next, we change the LSS component to a Gaussian distribution and observe how the results are affected, which can be done trivially with our pipeline. To that end, the SBI pipeline, as described in Section 4.2, is applied by changing the field in GLASS. For the comparison, we reduce the resolution of the simulations to $N_{\mathrm{side}}=512$. As we have established that the high-resolution run faithfully reproduces the real posterior distribution and is a good fit to the data in the last section, reducing the resolution to $N_{\mathrm{side}}=512$ makes a quantitative comparison while also requiring less computational resources. We also rerun the fiducial case at the same resolution as well to make a quantitative comparison. It should be noted that the lower resolution effectively reduces the field variance of the LSS component. In appendix A, we discuss the effect of reducing $N_{\mathrm{side}}$ on the coverage tests. It can be seen that the resulting posterior estimates are slightly biased when $N_{\mathrm{side}}$ is lowered, reducing the variance of the LSS component. Since we have established unbiased estimators at high resolution and checked that the constraints and $\chi^{2}$-test are not affected, this does not hamper our analysis when comparing different models.
@@
The contours are shown in Figure A.2 and the numerical values of the means and the $1\sigma$ errors of the parameters are presented in Table 2. As can be observed, there is a significant overlap among the values. Only by observing the contour plots we cannot distinguish the models. Hence, we rely on the $\chi^{2}$-tests, shown in Figure A.3, respectively. Even then, judging from these figures, there is no discernible difference between the Gaussian and log-normal LSS components. All the variations seem to be good fits according to the likelihood in Figure A.4. Consequently, with the current data, one cannot distinguish a Gaussian from a log-normal LSS component. A larger number of host-identified FRBs is required for meaningful detection of this difference. In particular, FRBs with higher redshift would be especially important to assess the effect due to the LSS field, as it becomes more dominant as the redshift increases. Following the simplified forecasts done in Hagstotz et al. (2022b), we would expect that a couple of hundred FRBs can distinguish between these different models.
@@
6.2.2 Variations in the Host component
Now, we turn our focus on the host model of the DM, changing it from the fiducial log-normal to first a truncated Gaussian (t-Gaussian hereafter) and then to a Gamma distribution. The Gaussian is truncated at zero as the DM is positive by definition, hence, it is not Gaussian mathematically. For the t-Gaussian and Gamma hosts, the prior on $\sigma$ is uniform on [0, 500], which is broad enough to capture the expected long-tail behaviour. It also implies that the Gamma distribution can resemble either a log-normal or a t-Gaussian distribution depending on the data, as all of these distributions come from the exponential family. As before, the contours, $\chi^{2}$ and likelihood are shown in Figure A.2 Figure A.3 and Figure A.4 respectively, all with $N_{\mathrm{side}}=512$.
@@ 6.2.2 Variations in the Host component
Looking at the best fit values, we can see that the statistical properties of the LSS field are rather sub-dominant with the current FRB sample and if a log-normal distribution is used for the host contribution. If we assume a (truncated) Gaussian host contribution, it cannot explain the high DM of FRBs at low redshifts, Hence, the model responds by artificially increasing the variance of the host contribution when the LSS is chosen to be Gaussian as well. In general, however, the constraints are all consistent with each other.
@@ 7 Conclusion
In the next step, we performed SBI on the described simulations. This inference method requires no explicit likelihood and works by training a NN to learn either the posterior distribution (non-amortised) or the joint distribution of data and parameters (amortised). In this context, the shape of the likelihood is unconstrained, which is why this approach is often referred to as likelihood-free inference (though it still implicitly requires a likelihood). To demonstrate our pipeline, we applied it to 12 host-identified FRBs as a function of the cosmological amplitude of the DM-$z$ relation, $\mathcal{A}$ in Equation 7, and two host contribution parameters. For our fiducial model, we used a log-normal electron density field as well as a log-normal host model. We inferred $\mathcal{A}=0.89^{+0.31}_{-0.29}$, consistent with unity or the Planck cosmology. Similarly, the median and scale of the host distribution are $\mathrm{DM}_{\mathrm{host}}=223^{+108}_{-95}\;\mathrm{pc\;cm}^{-3}\ \mathrm{and}\ \sigma_{\mathrm{LN}}=0.76^{+0.37}_{-0.44}$. The values for these three parameters are consistent with previous findings and we indeed found that no parameter (amplitude of the DM-$z$ relation, mean host contribution and its variance) is entirely dominated by its prior. This can be seen by the fact that the contours in Figure 3 are not entirely flat, even for $\sigma_{\mathrm{LN}}$. Following the discussion about the prior ranges in Table 1, this constraining power arises because very narrow ($\sigma_{\mathrm{LN}}\to 0$) and very broad ($\sigma_{\mathrm{LN}}>1.5$) are not supported by the data or by analytical models (see also Reischke et al., 2024). Therefore, there is additional information in the data. This indicates that already 12 FRBs can inform us about the shape of the host contribution to some extent.
@@
One of the added benefits of the SBI pipeline is that we can test the Gaussian likelihood assumption frequently used in a more traditional Bayesian approach. In that regard, we investigated the shape of the likelihood, finding the expected long-tailed distribution towards high DM values, which is necessary to explain the large DMs observed at low redshifts, such as those seen in FRB 20190520B. Another important check involved examining the evolution of the likelihood with redshift. We found that the mean of the probability distribution function increases with redshift, as expected. However, we also observed a reduction in its scatter. Although this behaviour may seem counterintuitive at first, it is supported by a straightforward analytical calculation, which shows that the width of the total distribution indeed reaches a minimum at redshift $z<1$ for the parameter values assumed in this analysis.
Lastly, we explored different modelling choices for both the LSS component and the host contribution. For the LSS component, we considered both Gaussian and log-normal distributions, while for the host contribution, we evaluated log-normal, truncated Gaussian and gamma distributions. We systematically tested all possible combinations of these models and found that they consistently provided similar constraints. To save computational time, we reduced the resolution of these simulations from $N_{\mathrm{side}}=4096$ to $N_{\mathrm{side}}=512$. We found that the TARP test indicates that the learned posterior in these cases is slightly biased. At lower $N_{\mathrm{side}}$, the variance of the LSS component is lower as small-scale effects are not accounted. Moreover, the interpolation of the HEALPix maps also smooths out pixel values due to low pixel counts at lower resolution. Combined, these two results in smaller variance, which lead to a biased coverage. However, as all models are compared with the same resolution, the results from the comparisons still hold. Importantly, none of the models showed any indication of being a poor fit to the data, which is mainly due to the low number of FRBs available with host identification at the moment.
In conclusion, the simulations and inference pipeline we have developed integrate the precise physical and statistical properties needed to accurately infer the DM-$z$ relation, free from general assumptions about the likelihood or posterior. This approach is highly adaptable and scales effectively with an increasing number of FRBs, thanks to the implemented data compression techniques. Moreover, our simulations provide a robust foundation for investigating systematic effects in cosmological studies involving FRBs, as these can be seamlessly incorporated at the map level. It is, for example, easy to introduce any residual MW component which was not entirely removed by the models. Furthermore, introducing redshift and localisation errors and unlocalised FRBs into the analysis can be done easily in the simulation; thus, looking ahead, a key direction will be to include more FRBs without known redshifts, enabling a joint fit of the DM-$z$ properties alongside the statistical properties of the DM. Implementing this in an analytic framework is intractable, making SBI an excellent alternative.
@@ Appendix A Effect of reducing $N_{\mathrm{side}}$
Figure A.1: The multivariate TARP coverages for the fiducial model (using a log-normal distribution for both the field and the host) These are computed from 1000 forward simulations. One can see that an increased resolution of the maps increases the accuracy of the final posterior estimate. This behaviour is found in all cases studied in this work.
In Figure A.1, we show the effect of reducing the resolution of the GLASS forward simulation on the TARP coverage. This is necessary to establish the baseline for the comparison of the different models for the host and the LSS contribution, which have been run at a lower resolution. The main effect of decreasing $N_{\mathrm{side}}$ is that power on small scales gets washed out. In particular, the maximum multipole properly resolved is $\ell_{\mathrm{max}}=3N_{\mathrm{side}}-1$ (Tessore et al., 2023). However, loss of power already occurs earlier. In terms of effects on the inference process, this means that the variance introduced by the LSS component decreases and therefore, the scatter in the data will be larger than in the simulations. Therefore, the final constraints might be artificially tight. Furthermore, inferring the DM contribution for a shell requires interpolation. Specifically, we take the bi-linear interpolation of the four neighbouring pixels. For a given $N_{\mathrm{side}}$, the total pixel count in a HEALPix map is $N_{\mathrm{pixel}}=12N^{2}_{\mathrm{side}}$ (Górski et al., 2005). Now, at a lower $N_{\mathrm{side}}$, the pixel count is lower and the interpolation effectively smooths out the expected fluctuation of $\mathrm{DM}_{\mathrm{LSS}}$, leading to an underestimate of error and bias in the TARP coverage. Figure A.1 demonstrates that the estimated posterior becomes slightly biased for decreasing $N_{\mathrm{side}}$ for the fiducial model. We also note that the bias is a combined effect of the two scenarios mentioned above. Since this will be true for all models and the fact that the high-resolution run is unbiased and exact, we can safely do a one-to-one comparison of the different models using a lower resolution.
One potential solution to reduce the bias is to use weighted sampling of the HEALPix maps to increase the LSS variance. However, we argue that it has little physical motivation. Therefore, as simulation-based inference (SBI) relies on the quality of simulations we use, it is advised to use a higher $N_{\mathrm{side}}$ for unbiased results. Alternatively, we can also reduce the threshold of the $1-\epsilon$ mass of the highest-probability region (HPR) of the approximate posterior in sequential training of TSNPE (more details can be found in Section 6.10 in Deistler et al., 2022).
Figure A.2: Marginalised contour plots for different combinations of the density field and host distribution. Each panel have the same host distribution and only the density field is varied, where the blue colour is for the log-normal field and red is for the Gaussian field. The host distribution for the top left, top right and bottom panels are log-normal, t-Gaussian and Gamma, respectively. The posteriors for all of these contours are evaluated with 1000 forward simulations in a 10-round TSNPE setup. The resolution of all the contours is $N_{\mathrm{side}}=512$ for comparison. The means with the $1\sigma$ errors are provided in Table 2.
@@ Appendix B Results from different fits
In this section, we present a comprehensive analysis of the different figures resulting from variations in both the host contribution and the large-scale structure (LSS) contribution. Specifically, we display the final contour plots Figure A.2, the goodness-of-fit tests in Figure A.3 and the likelihoods in Figure A.4. Note that we use $N_{\mathrm{side}}=512$ in this comparison to conserve computational resources. The reduced resolution underestimates the error from cosmic variance on individual measurements, leading to artificially tighter constraints. This issue has been addressed in our fiducial run, where a higher $N_{\mathrm{side}}$ was utilised to validate the pipeline, as described in the main text. The key conclusion from these analyses is that all combinations of the model are statistically consistent with one another and provide a satisfactory fit to the data. This is primarily due to the relatively small number of FRBs considered in this study, as well as the conservative error estimates and the expansive posterior volume resulting from the use of a three-parameter model. While these effects limit the current discriminative power, they ensure robustness in our findings.
@@
Finally, for reference, we provide a list of the FRBs used in this analysis in Footnote 3.
@@
Figure A.4: Likelihoods for all the models in our analysis. The two rows represent the log-normal and Gaussian density fields and the three columns are for the log-normal, t-Gaussian and Gamma host distribution. The simulated DMs for the individual FRBs are shown with a colour bar by mapping them to the percentile values. The cyan dots are the median and the black crosses are the observed DM.
@@
Appendix C Validation with mock catalogue
To maintain end-to-end functionality and pipeline reliability, we ran the full analysis using a mock catalogue. The inputs in our pipeline are the right ascension (RA), declination (DEC), redshift ($z$), the Milky Way contribution ($\mathrm{DM}_{\mathrm{MW}}$) and observed dispersion measure ($\mathrm{DM}_{\mathrm{obs}}$) of the FRBs. For the test, we create a mock catalogue of 100 FRBs, where the RA and DEC are random samples drawn from a uniform distribution. Choosing the appropriate redshift distribution of FRBs is non-trivial as it closely relates to the formation mechanism, an area of ambiguity and active research. The distribution alters depending on the model, i.e. tracking the star formation history or compact mergers (see e.g. Muñoz et al., 2016; Wang et al., 2020; Reischke et al., 2021; James et al., 2022a; Peng & Yu, 2025). Hence, we use a more general distribution where the PDF is $z^{2}\mathrm{exp}(-3z)$ and peaks at $z\approx 1$. Additionally, we consider a redshift range of $z=[0.1,1]$, which mimics our analysis with the real data. The values of $\mathrm{DM}_{\mathrm{MW}}$ are taken from the YMW16 model (Yao et al., 2017). Finally, the simulator is run to create the mock dispersion ($\mathrm{DM}_{\mathrm{obs}}$). We use the fiducial model with a log-normal LSS field and log-normal host configuration where the parameters are set as $\mathcal{A}=1,\mathrm{DM}_{\mathrm{host}}=200$ and $\sigma_{\mathrm{LN}}=0.35$. This constitutes the true values for the mocks, which are then used for the inference.
In Figure A.5, we show the contours from that analysis. The resolution parameter is $N_{\mathrm{side}}=512$ and all other parameters remain unaltered from the principal analysis. The mean values, indicated by the black line, are $\mathcal{A}=0.96\pm 0.11,\mathrm{DM}_{\mathrm{host}}=284^{+100}_{-200}\;\mathrm{pc\;\mathrm{cm}^{-3}}$ and $\sigma_{\mathrm{LN}}=0.54^{+0.22}_{-0.44}$. It demonstrates that the analysis can recover the true values (red line) effectively as the true values lie well within the $1\sigma$ range. A noteworthy aspect is that, despite having many more FRBs, the mock FRB sample is at higher redshifts than the real FRB sample. Therefore the host contribution is less dominant and therefore less constrainted than in the real analysis presented in this work. In contrast, however, the amplitude is much better constrained in the mock analysis for the same reason. Moreover, we check the coverage of the contours with the TARP coverage test and the result is shown in Figure A.6.
Figure A.5: Marginalised contour plots with 100 mock FRB samples. The fiducial model with log-normal LSS field and log-normal host distribution is used. The truth (red line) is at (1, 200, 0.35) while the mean (black line) is at $\mathcal{A}=0.96\pm 0.11,\mathrm{DM}_{\mathrm{host}}=284^{+100}_{-200}\;\mathrm{pc\;\mathrm{cm}^{-3}}$ and $\sigma_{\mathrm{LN}}=0.54^{+0.22}_{-0.44}$.
Figure A.6: TARP coverage for the contour in Figure A.5. The simulations are with $N_{\mathrm{side}}=512$.